80 = -31.08 + 22.26ln(x)

How do I solve this equation for x?
Answer #1

Bo 803-31,08 + 22.26 In 1) (Given equation -31.08+22.26 In (x) =80 (Rewritin 22.2616 (2) = 80+31.08 22,26101 (C) = 11:08 (Rew